OpenStudy (gelow):

What is the perimeter of a triangle with the given side lengths? s1 = x + 4 cm s2 = 3x + 1 cm s3 = 7x + 3 cm P = x + cm

OpenStudy (gelow):

@KamiBug

OpenStudy (gelow):

@Crissy15 @Arianna842

OpenStudy (gelow):

@HelpBlahBlahBlah

OpenStudy (anonymous):

S1 + s2+s3 = (x+4cm) +(3x+1cm)+(7x+3cm) = (x + 3x + 7x) + (4cm + 1cm + 3cm) = 11x + 8cm P=11x + 8cm

OpenStudy (gelow):

ok now i get it thx you got a medal : )

OpenStudy (anonymous):

thankyou!